OpenStudy (anonymous):

tan(sin^-1)(-1)= a. 0 b. -1 c. undefined d. 1

OpenStudy (anonymous):

\[\theta = \sin^{-1} \left( -1 \right)\]\[\sin \theta = -1\]\[\theta = 270^{\circ}\]\[\tan \left( \sin^{-1} \left( -1 \right) \right) = \tan \theta = \tan 270^{\circ} = undefined\]